How many gallons of 60% antifreeze solution must be mixed with 100 gallons of 20% antifreeze to get a mixture that is 50% antifreeze? Use the six-step method.
Answer by mananth(16946)

-------------- percent ---------------- Amount
Antifreeze I 20 ---------------- 100 gal
Antifreeze II 60 ---------------- x gal
Mixture 50 ---------------- 100 + x gal

20 * 100 + 60 x = 50 ( 100 + x )
2000 + 60 x = 5000 + 50 x
60 x -50 x = -2000 + 5000
10 x = 3000
/ 10
x = 300 gal Antifreeze II

m.ananth@hotmail.ca